If 12×7=96 and 9×9=36 then 13×4=?tell me the answer

Grade:6

2 Answers

Yukti
6 years ago
12×7=96 9×9=3612=96/7 ---¡ 9=36/9So 13 = 12+1 =4 From ¡ therefore subtracting 4 13=96/7+1 from 9 =103/7 9-5 =4-5=-1Hence 13×4=103/7×-1=-103/7
Vivek
6 years ago
Answer =26 there is pattern that that 1st no. Is multiplied by 8 in 1st expresion and in 2nd its by 4 now in third it must be by 2 so 13×2= 26.